OpenStudy (anonymous):

The top of a gazebo is to be painted at a cost of $0.14 per square yard. The rectangular part of the gazebo is 125 yards long and the diameter of each semicircle is 47 yards. Find the cost of painting the gazebo. Use 3.14 for p.

OpenStudy (anonymous):

I have no idea what kind of gazebo has rectangles and semicircles. Is there a diagram you didn't show us? We need to see it.

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

Did the image of the diagram go through

OpenStudy (radar):

That is one BIG gazebo.

OpenStudy (anonymous):

do you know were to put the numbers

OpenStudy (anonymous):

No.

OpenStudy (radar):

Larger than a foot-ball field!

OpenStudy (anonymous):

Let's not worry about the cost first, and do the math for the area of the gazebo. Okay, so what do two semicircles equal? One circle. Now we can treat the two semicircles as a circle so we have one circle and one rectangle. The rectangle part is easy. 125 yards long * 47 yards. Where'd we get the 47 from? Well, the DIAMETER of the circle is what it measures across, which is equal to the width of the rectangle in the diagram. 125*47 = 5875 square yards Now we've got the first part. Let's get the area of the circle. The area of a circle is pi times it's radius squared. How do we get the radius? Well, the radius is just half the diameter. 47/2 is the radius, and pi is 3.14 (47/2)^2 * 3.14 = 1734.065 So now we add it all together, 1734 + 5875 = 7609 square yards. Let's multiply it by 0.14 now, and we get $1065.26

OpenStudy (radar):

Calculate the area of the rectangular portion (125 by 47) yards Calculate the area of a circle with a radius of 23.5 yds that will take care of the two semicircles, Add.

OpenStudy (radar):

I am amazed it could be painted at 14 cents a square yd.......that wouldn't even pay for the paint.